# Gibbs-preserving operations requiring infinite amount of quantum coherence

### v2

@@ Appendix C Construction of pairwise reversible Gibbs-preserving Operations
Appendix D State transformations requiring cost-diverging Gibbs preserving Operations
(Proof of Theorem 4)
@@ Appendix D State transformations requiring cost-diverging Gibbs preserving Operations
Besides the physical implementability of a given Gibbs-preserving Operation, another significant problem is whether a given state transformation $\rho\to\sigma$ that can be realized by a Gibbs-preserving Operation can physically be achieved by a finite coherence cost. Theorem 2 or 3 does not directly address this question because there are typically many quantum channels that can realize the said state transformation. To investigate the smallest coherence cost for a given state transformation, one needs to optimize over all such quantum channels, which is generally intractable.
@@
To this end, we focus on the setting where an operation implemented by a Thermal Operation with coherence assistance is also Gibbs-preserving. This is a reasonable restriction given that we are considering a state transformation that can be realized by a Gibbs-preserving Operation. Here, we employ our methods to show that some state transformations realized by Gibbs-preserving Operations can never be achieved by a Thermal Operation and finite coherence, under the restriction that the implemented opearation itself is Gibbs-preserving.
Theorem S.6 (Theorem 4 in the main text).
@@
$\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j},$ (67)
there is a pair $(\rho,\sigma)$ of states on $S$ and $S^{\prime}$ satisfying the following two conditions:
  • $\rho\rightarrow\sigma$ is possible by a Gibbs-preserving Operation from $S$ to $S^{\prime}$.
  • There is no Gibbs-preserving Operation $\Lambda$ such that $\Lambda(\rho)=\sigma$ and ${\cal F}_{c}^{\epsilon=0}(\Lambda)$ is finite.
@@
$|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}}$ (68)
@@
$|\eta_{+}\rangle_{S}$ $:=\sqrt{r}|i\rangle_{S}+\sqrt{1-r}|j\rangle_{S},$ (69)
$r$ $:=\frac{\tau_{S^{\prime},i^{\prime}}-\tau_{S,j}}{\tau_{S,i}-\tau_{S,j}}.$ (70)
We first show that there exists a Gibbs-preserving Operation that realizes the transformation (68). Consider a map $\Lambda_{0}:S\to S^{\prime}$ defined by
$\Lambda_{0}(...)$ $:=\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|...]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}+\mathrm{Tr}[(1-|\eta_{+}\rangle\langle\eta_{+}|)...]\xi_{S^{\prime}},$ (71)
$\xi_{S^{\prime}}$ $:=\frac{\tau_{S^{\prime}}-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}}{1-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]}.$ (72)
@@
It now suffices to show that an arbitrary Gibbs-preserving Operation $\Lambda:S\to S^{\prime}$ that realizes the transformation (68) must come with diverging coherence cost, i.e., ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$. Let $\{M_{k}\}_{k}$ be a Kraus representation of $\Lambda$. We write each Kraus operator with respect to energy eigenbases $\{|l\rangle_{S}\}_{l=1}^{d_{S}}$ of $S$ and $\{|m\rangle_{S^{\prime}}\}_{m=1}^{d_{S^{\prime}}}$ of $S^{\prime}$ as
@@
$+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{l\neq i,j}e_{k,l}|i^{\prime}\rangle_{S^{\prime}}\langle l|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S},$ (73)
where $|\eta_{-}\rangle:=\sqrt{1-r}|j\rangle_{S}-\sqrt{r}|i\rangle_{S}$.
@@
$b_{k,m}=0,\enskip\forall k,\enskip\forall m\mbox{ s.t. }m\neq i^{\prime}.$ (74)
@@
$\sum_{k}|a_{k}|^{2}$ $=\sum_{k}\langle\eta_{+}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=1,$ (75)
$\sum_{k}c^{*}_{k}a_{k}$ $=\sum_{k}\langle\eta_{-}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,$ (76)
$\sum_{k}e^{*}_{k,l}a_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j$ (77)
$\sum_{k}e^{*}_{k,l}c_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{-}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (78)
@@
$c_{k}$ $=0,\enskip\forall k$ (80)
$e_{k,l}$ $=0,\enskip\forall k,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (81)
@@
$M_{k}=$ $a_{k}|i^{\prime}\rangle_{S^{\prime}}\langle\eta_{+}|_{S}+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S}.$ (82)
@@
$\langle i^{\prime}|_{S^{\prime}}\Lambda(\psi_{S})|i^{\prime}\rangle_{S^{\prime}}=0,\enskip\forall|\psi\rangle_{S}\mbox{ s.t. }\langle{\eta_{+}}|{\psi}\rangle_{S}=0.$ (83)

### v3

@@ III Construction of pairwise reversible Gibbs-preserving Operations (Proof of Theorem 2)
IV Upper bound for the coherence cost of the channel given in Eq. (2)
$\Lambda(\rho)=\langle{1}|{\rho}|{1}\rangle\eta+\langle{0}|{\rho}|{0}\rangle\sigma$ (S.89)
satisfies the upper bound
${\cal F}_{c}(\Lambda)\leq{\cal F}(\eta)+{\cal F}(\sigma).$ (S.90)
Proof.
$\Lambda_{1}(\kappa_{S}):=\langle 1|\kappa_{S}|1\rangle|{1}\rangle\!\langle{1}|_{E}+\langle 0|\kappa_{S}|0\rangle|{0}\rangle\!\langle{0}|_{E}$ (S.91)
and
$\Lambda_{2}(\kappa_{E}):=\langle 1|\kappa_{E}|1\rangle\eta+\langle 0|\kappa_{E}|0\rangle\sigma$ (S.92)
for arbitrary states $\kappa_{S}$ in $S$ and $\kappa_{E}$ in $E$.
$\Lambda_{1}(\kappa_{S})$ $=\mathrm{Tr}_{S}[U\kappa_{S}\otimes|0\rangle\langle 0|_{E}U^{\dagger}]$ (S.93)
$\Lambda_{2}(\kappa_{E})$ $=\mathrm{Tr}_{E\tilde{S}}[V\kappa_{E}\otimes\eta\otimes\sigma V^{\dagger}]$ (S.94)
where
$U$ $:=|00\rangle\langle 00|_{SE}+|11\rangle\langle 10|_{SE}+|01\rangle\langle 01|_{SE}+|10\rangle\langle 11|_{SE},$ (S.95)
$V$ $:=|0\rangle\langle 0|_{E}\otimes\mathds{1}_{S\tilde{S}}+|1\rangle\langle 1|_{E}\otimes U_{\mathrm{SWAP}}$ (S.96)
${\cal F}_{c}(\Lambda_{1})$ $\leq{\cal F}(|{0}\rangle\!\langle{0}|)=0$ (S.97)
${\cal F}_{c}(\Lambda_{2})$ $\leq{\cal F}(\eta\otimes\sigma)={\cal F}(\eta)+{\cal F}(\sigma).$ (S.98)
Since ${\cal F}_{c}(\Lambda)\leq{\cal F}_{c}(\Lambda_{1})+{\cal F}_{c}(\Lambda_{2})$, we obtain (S.90). ∎
We note that (S.90) is only an upper bound of the cost of the channel in (2) of the main text, which may not be tight in general.
V On atypicality of cost-diverging Gibbs-preserving Operations
Theorem S.2 and Corollary S.7 show that there are infinitely many cost-diverging Gibbs-preserving Operations. However, this does not imply that the ones with diverging coherence cost are typical instances among all Gibbs-preserving Operations. Here, we show that, in fact, cost-diverging operations are measure zero with respect to the set of all Gibbs-preserving Operations. Nevertheless, we stress that one still needs to be cautious about the implementability of Gibbs-preserving Operations because (1) in many cases, we are interested in highly structured Gibbs-preserving Operations, which can be cost-diverging ones, and (2) even if a desired Gibbs-preserving Operation has a finite coherence cost, it can still come with a large coherence cost. Indeed, Theorem S.2 implies that the operations that approximate a cost-diverging one with accuracy $\epsilon$ must come with a coherence cost proportional to $1/\epsilon$, which forms a set with nonzero measure.
We show the atypicality of cost-diverging instances by proving that they only reside on the boundary of the set of Gibbs-preserving Operations. To formalize this, we need the notion of the interior points of a subset of vector spaces, where the subset of interest itself could also have zero measure. This is represented by the relative interior of the subset Rockafellar (2015). For a vector space equipped with trace norm, define the open ball centered around a vector $v$ by
[…]
Also, we define the intersection between the open ball and the affine hull of a set ${\cal S}$ by
${\cal B}^{\epsilon}_{\cal S}(v)\coloneqq{\cal B}^{\epsilon}(v)\cap{\rm aff}({\cal S})$ (S.100)
[…]
In the following, we show that every operation in the relative interior of the set of Gibbs-preserving Operations has a finite coherence cost, which implies that the cost-diverging ones are measure zero with respect to all Gibbs-preserving Operations because any finite-dimensional convex set has a nonzero relative interior (Rockafellar, 2015, Theorem 6.2). To this end, we begin by the following result, which we will employ to show that if all members in the set can be approximated arbitrarily well, all interior points of this set can be exactly implemented. This was shown in Ref. (Wilming, 2022, Corollary 7) for interior points of the full-dimensional subset, and here we extend it to relative interior points. Nevertheless, the proof is almost identical, and here we particularly follow the descriptions employed in Ref. (Shiraishi and Takagi, 2024, Lemma S.14), which explained the proof in a bit more detail.
Lemma S.9.
Let $\mathcal{V}$ be a subset of linear operators on a finite-dimensional Hilbert space, and let $\{\mathcal{S}_{n}\}_{n=1}^{\infty}$ be a family of closed convex sets satisfying $\mathcal{S}_{n}\subseteq\mathcal{V}$ and $\mathcal{S}_{n}\subseteq\mathcal{S}_{n+1}$ for every $n$. Also, suppose that $\{\mathcal{S}_{n}\}_{n=1}^{\infty}$ approximates $\mathcal{V}$ arbitrarily well, in the sense that for an arbitrary $\epsilon>0$, there exists a sufficiently large $N$ such that $\min_{u\in\mathcal{S}_{N}}\|v-u\|_{1}<\epsilon$ for every $v\in\mathcal{V}$. Then, for every $\kappa\in{\rm relint}(\mathcal{V})$, there exists an integer $n$ such that $\kappa\in\mathcal{S}_{n}$.
Proof.
Since $\kappa\in{\rm relint}(\mathcal{V})$, there exists a sufficiently small $\delta>0$ such that $B_{\mathcal{V}}^{\delta}(\kappa)\subseteq\mathcal{V}$. By assumption of $\{\mathcal{S}_{n}\}_{n}$, there exists an integer $N_{\delta/2}$ such that $\min_{u\in\mathcal{S}_{N_{\delta/2}}}\|v-u\|_{1}<\delta/2$ for an arbitrary $v\in\mathcal{V}$.
In the following, we aim to show that $\kappa\in\mathcal{S}_{N_{\delta/2}}$. Suppose contrarily that $\kappa\not\in\mathcal{S}_{N_{\delta/2}}$. Let $s\in\mathcal{S}_{N_{\delta/2}}$ be the closest point in $\mathcal{S}_{N_{\delta/2}}$ from $\kappa$, i.e.,
$s={\rm argmin}_{s^{\prime}\in\mathcal{S}_{N_{\delta/2}}}\|\kappa-s^{\prime}\|_{1}.$ (S.102)
Letting $r(t)=(1-t)s+t\kappa$, it holds that for every $\delta^{\prime}<\delta$ there exists $t^{\prime}\geq 1$ such that $\|r(t^{\prime})-\kappa\|_{1}=(t^{\prime}-1)\|s-\kappa\|_{1}=\delta^{\prime}$. Together with the fact that $r(t)\in{\rm aff}(\mathcal{V})$, which is because $s\in\mathcal{S}_{N_{\delta/2}}\subseteq\mathcal{V}$ and $\kappa\in\mathcal{V}$, we have $r(t^{\prime})\in{\cal B}^{\delta}_{\mathcal{V}}(\kappa)\subseteq\mathcal{V}$. Choosing $\delta^{\prime}=2\delta/3$ in particular, we get
$\|r(t^{\prime})-s\|_{1}=t^{\prime}\|s-\kappa\|_{1}\geq(t^{\prime}-1)\|s-\kappa\|_{1}=\|r(t^{\prime})-\kappa\|_{1}=2\delta/3.$ (S.103)
On the other hand,
$\min_{u\in\mathcal{S}_{N_{\delta/2}}}\|r(t^{\prime})-u\|_{1}$ $=\min_{u\in\mathcal{S}_{N_{\delta/2}}}\|-(t^{\prime}-1)s+t^{\prime}\kappa-u\|_{1}$ (S.104)
$=t^{\prime}\min_{u\in\mathcal{S}_{N_{\delta/2}}}\left\|\kappa-\left[\left(1-\frac{1}{t^{\prime}}\right)s+\frac{1}{t^{\prime}}u\right]\right\|_{1}=t^{\prime}\|\kappa-s\|_{1}$
where in the last equality, we used the fact that $(1-\frac{1}{t^{\prime}})s+\frac{1}{t^{\prime}}u\in\mathcal{S}_{N_{\delta/2}}$ because of the convexity of $\mathcal{S}_{N_{\delta/2}}$ and (S.102), ensuring $u=s$ achieves the minimum. Combing (S.103) and (S.104), we get
$\min_{u\in\mathcal{S}_{N_{\delta/2}}}\|r(t^{\prime})-u\|_{1}\geq 2\delta/3.$ (S.105)
Since the definition of $\mathcal{S}_{N_{\delta/2}}$ implies $\delta/2\geq\min_{u\in\mathcal{S}_{N_{\delta/2}}}\|r(t^{\prime})-u\|_{1}$, we reach a contradiction $\delta/2\geq 2\delta/3$ together with (S.105). ∎
The next result shows that if a subset contains an interior point of its superset, no interior points of the subset cannot be on the boundary of the superset. We will employ this to ensure that relative interior of the set of Gibbs-preserving Operations are also in the interior of all quantum channels.
Lemma S.10.
Let ${\cal S}\subseteq{\cal V}$ be a subset that contains at least one interior point $v\in{\rm relint}({\cal V})$ with respect to ${\cal V}$. Then, every $s\in{\rm relint}({\cal S})$ is also a relative interior point of ${\cal V}$, i.e., $s\in{\rm relint}({\cal V})$.
Proof.
For an arbitrary $s\in{\rm relint}({\cal S})$, there exists $\epsilon>0$ such that ${\cal B}_{\cal V}^{\epsilon}(s)\subseteq{\cal V}$. For $v\in{\rm relint}({\cal V})$, let $r(t)=-tv+(1+t)s$ for $t\geq 0$. Since $r(t)\in{\rm aff}({\cal S}),\forall t$ and $s\in{\rm relint}({\cal S})$, there exists sufficiently small $\tilde{t}>0$ such that $r(\tilde{t})\in{\cal S}\subseteq{\cal V}$. Recalling the line segment property (Rockafellar, 2015, Theorem 6.1), which states that for an arbitrary closed convex set ${\cal C}$ and arbitrary points $x\in{\rm relint}({\cal C})$ and $y\in{\cal C}$, all points $\lambda x+(1-\lambda)y$ with $0<\lambda\leq 1$ are in the relative interior of ${\cal C}$, $s=\frac{r(\tilde{t})+\tilde{t}v}{1+\tilde{t}}$ satisfies $s\in{\rm relint}({\cal V})$ as $v\in{\rm relint}(\cal V)$ and $r(\tilde{t})\in{\cal V}$, as well as $\tilde{t}>0$. ∎
We are in the position to show the desired statement by combining Lemmas S.9 and S.10, together with the fact that every quantum channel can be approximated arbitrarily well with a finite coherence cost ensured by Theorem S.13.
Proposition S.11.
Let $\mathbb{O}_{\rm GPO}$ be the set of Gibbs-preserving Operations. Then, every coherence-diverging Gibbs-preserving Operation $\Lambda$ must be on the boundary of $\mathbb{O}_{\rm GPO}$, i.e., $\Lambda\not\in{\rm relint}(\mathbb{O}_{\rm GPO})$.
Proof.
We show that every $\Lambda\in{\rm relint}(\mathbb{O}_{\rm GPO})$ can be implemented exactly with a finite coherence cost. For a subset $\mathbb{O}$ of channels, let $\tilde{\mathbb{O}}$ denote the set of Choi operators for $\mathbb{O}$. Let $\mathbb{O}_{\rm GPO}(S\to S^{\prime})$ be the set of Gibbs-preserving Opeations from $S$ to $S^{\prime}$, and let $\mathbb{T}(S\to S^{\prime})$ be the set of all channels from $S$ to $S^{\prime}$.
[…]
Let $\mathcal{S}_{n}$ be the set of Choi operators of channels with input and output systems $S$ and $S^{\prime}$ whose coherence cost is upper bounded by the integer $n$, i.e.,
[…]
where ${\rm cl}$ refers to the closure of the set, whose necessity comes from that in the implementation $\Xi(\cdot)=\Lambda(\cdot\otimes\eta),\Lambda\in\mathbb{O}_{\rm TO}$, the choice of the system that $\eta$ acts on is arbitrary and thus the set of states $\eta$ that admits ${\cal F}_{c}^{\epsilon=0}(\Xi)\leq n$ may not be closed.
We can see that $\mathcal{S}_{n}$ is convex, following a similar argument to the one in (Lostaglio et al., 2015, Appendix C). Consider $\Xi_{1},\Xi_{2}\in{\cal S}_{n}$ that admit implementation $\Xi_{i}(\cdot)=\mathrm{Tr}_{\overline{S}^{\prime}}[U_{i}(\cdot\otimes\eta_{i}\otimes\tau_{i})U_{i}^{\dagger}],i=1,2$ where $U_{i}$ is an energy-conserving unitary on $SR_{i}E_{i}$, $\tau_{i}$ is the thermal state of some system $E_{i}$, and $\eta_{i}$ is a coherent resource state on some system $R_{i}$ with ${\cal F}(\eta_{i})\leq n$. Note that $\mathrm{Tr}_{\overline{S}^{\prime}}$ denotes the partial trace over systems other than $S^{\prime}$. It then suffices to show that $\Xi=p\Xi_{1}+(1-p)\Xi_{2}$ can be implemented by a Thermal Operation aided by a resource state $\eta$ with ${\cal F}(\eta)\leq n$. To this end, let $\eta=p|{0}\rangle\!\langle{0}|_{F}\otimes\tilde{\eta}_{1}+(1-p)|{1}\rangle\!\langle{1}|_{F}\otimes\tilde{\eta}_{2}$ where $\tilde{\eta}_{1}=\eta_{1}\oplus 0_{R_{2}}$ and $\tilde{\eta}_{2}=0_{R_{1}}\oplus\eta_{2}$ are states on $R=R_{1}\oplus R_{2}$ with Hamiltonian $H_{R}=H_{R_{1}}\oplus H_{R_{2}}$, and $F$ is the system introduced for a flag with trivial Hamitlonian $H_{F}=0$. Let $U=|{0}\rangle\!\langle{0}|_{F}\otimes\tilde{U}_{1}\otimes\mathbb{I}_{E_{2}}+|{1}\rangle\!\langle{1}|_{F}\otimes\tilde{U}_{2}\otimes\mathbb{I}_{E_{1}}$ where $\tilde{U}_{1}=U_{1}\oplus\mathbb{I}_{SR_{2}E_{1}}$ and $\tilde{U}_{2}=U_{2}\oplus\mathbb{I}_{SR_{1}E_{2}}$. This is energy-conserving because
$[U,H_{F}+H_{S}+H_{R}+H_{E_{1}}+H_{E_{2}}]$ (S.107)
$=|{0}\rangle\!\langle{0}|\otimes[U_{1},H_{S}+H_{R_{1}}+H_{E_{1}}]+|{1}\rangle\!\langle{1}|\otimes[U_{2},H_{S}+H_{R_{2}}+H_{E_{2}}]=0.$
Then,
$\mathrm{Tr}_{\overline{S}^{\prime}}[U(\rho\otimes\eta\otimes\tau_{1}\otimes\tau_{2})U^{\dagger}]$ $=p\mathrm{Tr}_{\overline{S}^{\prime}}[U_{1}(\rho\otimes\eta_{1}\otimes\tau_{1})U_{1}^{\dagger}]+(1-p)\mathrm{Tr}_{\overline{S}^{\prime}}[U_{2}\rho\otimes\eta_{2}\otimes\tau_{2}U_{2}^{\dagger}]$ (S.108)
$=p\Xi_{1}+(1-p)\Xi_{2}.$
Noting that ${\cal F}(\eta)=p{\cal F}(\eta_{1})+(1-p){\cal F}(\eta_{2})\leq n$ concludes the proof that ${\cal S}_{n}$ is convex.
Theorem S.13 reveals that every quantum channel can be implemented with an arbitrary accuracy $\epsilon$ with a finite coherence cost specified in the right-hand side of (S.156). This particularly means that every Choi operator in $\tilde{\mathbb{T}}(S\to S^{\prime})$ can be approximated with arbitrary accuracy by a Choi operator in ${\cal S}_{N}\subseteq\tilde{\mathbb{T}}(S\to S^{\prime})$ for a sufficiently large $N$. Therefore, applying Lemma S.9 with $\mathcal{V}=\tilde{\mathbb{T}}(S\to S^{\prime})$, we get that for every $J_{\Lambda}\in{\rm relint}(\tilde{\mathbb{O}}_{\rm GPO}(S\to S^{\prime}))\subseteq{\rm relint}(\tilde{\mathbb{T}}(S\to S^{\prime}))$, where the second inclusion is guaranteed by Lemma S.10, and for every $\delta>0$ such that ${\cal B}_{\cal V}^{\delta}(J_{\Lambda})\subseteq{\rm relint}({\cal V})$, there exists an integer $n$ such that ${\cal B}_{\cal V}^{\delta}(J_{\Lambda})\subseteq{\cal S}_{n}$. This particularly means that $J_{\Lambda}\in{\rm relint}({\cal S}_{n})$ and thus ${\cal F}_{c}^{\epsilon=0}(\Lambda)\leq n$, i.e., $\Lambda$ can be implemented exactly with a finite coherence cost.
@@ V On atypicality of cost-diverging Gibbs-preserving Operations
VI State transformations requiring cost-diverging Gibbs preserving Operations (Proof of Theorem 3)
@@
Besides the physical implementability of a given Gibbs-preserving Operation, another significant problem is whether a given state transformation $\rho\to\sigma$ that can be realized by a Gibbs-preserving Operation can physically be achieved by a finite coherence cost. Theorem 1 or 2 does not directly address this question because there are typically many quantum channels that can realize the said state transformation. To investigate the smallest coherence cost for a given state transformation, one needs to optimize over all such quantum channels, which is generally intractable.
@@
To this end, we focus on the setting where an operation implemented by a Thermal Operation with coherence assistance is also Gibbs-preserving. This is a reasonable restriction given that we are considering a state transformation that can be realized by a Gibbs-preserving Operation. Here, we employ our methods to show that some state transformations realized by Gibbs-preserving Operations can never be achieved by a Thermal Operation and finite coherence, under the restriction that the implemented opearation itself is Gibbs-preserving. We also show that the cost-diverging state transitions given by our method are costly even when we allow small errors in their realization. To describe the result, we define the cost of approximate state transitions as follows:
[…]
Here, $\mathbb{O}_{\rm GPO}$ is the set of Gibbs-preserving Operations.
Theorem S.12 ( A generalization of Theorem 3 in the main text).
@@
$\tau_{S,i}<\tau_{S^{\prime},i^{\prime}}<\tau_{S,j},$ (S.110)
there is a pair $(\rho,\sigma)$ of states on $S$ and $S^{\prime}$ satisfying (i)$\rho\rightarrow\sigma$ is possible by a Gibbs-preserving Operation from $S$ to $S^{\prime}$, (ii) $\rho\rightarrow\sigma$ is impossible by any finite-cost Gibbs-preserving Operations from $S$ to $S^{\prime}$, i.e. ${\cal F}^{\epsilon=0}(\rho\rightarrow\sigma)=\infty$, and (iii) for non-zero error $\epsilon$, $\rho\rightarrow\sigma$ satisfies
${\cal F}^{\epsilon}(\rho\rightarrow\sigma)\geq\frac{\kappa(\epsilon,\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})}{\epsilon}-(\Delta(H_{S})+3\Delta(H_{S^{\prime}}))^{2},$ (S.111)
where $\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j})$ is the following real-valued function that is positive when (S.110) holds:
$\kappa(\beta,\tau_{S,i},\tau_{S^{\prime},i^{\prime}},\tau_{S,j}):=\frac{(\tau_{S,i}+\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})\log\frac{\tau_{S,j}}{\tau_{S,i}}}{2\beta(\tau_{S,j}-\tau_{S,i})^{2}(\tau_{S,i}+\tau_{S,j}+2\sqrt{(\tau_{S,j}-\tau_{S^{\prime},i^{\prime}})(\tau_{S^{\prime},i^{\prime}}-\tau_{S,i})})}$ (S.112)
We remark that (i) and (ii) correspond to the Theorem 3 in the main text.
@@
$|\eta_{+}\rangle_{S}\langle\eta_{+}|_{S}\rightarrow|i^{\prime}\rangle_{S^{\prime}}\langle i^{\prime}|_{S^{\prime}}$ (S.113)
@@
$|\eta_{+}\rangle_{S}$ $:=\sqrt{r}|i\rangle_{S}+\sqrt{1-r}|j\rangle_{S},$ (S.114)
$r$ $:=\frac{\tau_{S^{\prime},i^{\prime}}-\tau_{S,j}}{\tau_{S,i}-\tau_{S,j}}.$ (S.115)
We first show that there exists a Gibbs-preserving Operation that realizes the transformation (S.113). Consider a map $\Lambda_{0}:S\to S^{\prime}$ defined by
$\Lambda_{0}(...)$ $:=\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|...]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}+\mathrm{Tr}[(1-|\eta_{+}\rangle\langle\eta_{+}|)...]\xi_{S^{\prime}},$ (S.116)
$\xi_{S^{\prime}}$ $:=\frac{\tau_{S^{\prime}}-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]|i^{\prime}\rangle\langle i^{\prime}|_{S^{\prime}}}{1-\mathrm{Tr}[|\eta_{+}\rangle\langle\eta_{+}|\tau_{S}]}.$ (S.117)
@@
Next, we show (ii), i.e., the state transition cannot be achieved by finite-cost Gibbs-preserving operations. It suffices to show that an arbitrary Gibbs-preserving Operation $\Lambda:S\to S^{\prime}$ that realizes the transformation (S.113) must come with diverging coherence cost, i.e., ${\cal F}_{c}^{\epsilon=0}(\Lambda)=\infty$. Let $\{M_{k}\}_{k}$ be a Kraus representation of $\Lambda$. We write each Kraus operator with respect to energy eigenbases $\{|l\rangle_{S}\}_{l=1}^{d_{S}}$ of $S$ and $\{|m\rangle_{S^{\prime}}\}_{m=1}^{d_{S^{\prime}}}$ of $S^{\prime}$ as
@@
$+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{l\neq i,j}e_{k,l}|i^{\prime}\rangle_{S^{\prime}}\langle l|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S},$ (S.118)
where $|\eta_{-}\rangle:=\sqrt{1-r}|i\rangle_{S}-\sqrt{r}|j\rangle_{S}$.
@@
$b_{k,m}=0,\enskip\forall k,\enskip\forall m\mbox{ s.t. }m\neq i^{\prime}.$ (S.119)
@@
$\sum_{k}|a_{k}|^{2}$ $=\sum_{k}\langle\eta_{+}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=1,$ (S.120)
$\sum_{k}c^{*}_{k}a_{k}$ $=\sum_{k}\langle\eta_{-}|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,$ (S.121)
$\sum_{k}e^{*}_{k,l}a_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{+}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j$ (S.122)
$\sum_{k}e^{*}_{k,l}c_{k}$ $=\sum_{k}\langle l|M^{\dagger}_{k}M_{k}|\eta_{-}\rangle=0,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (S.123)
@@
$c_{k}$ $=0,\enskip\forall k$ (S.125)
$e_{k,l}$ $=0,\enskip\forall k,\enskip\forall l\mbox{ s.t. }l\neq i,j.$ (S.126)
@@
$M_{k}=$ $a_{k}|i^{\prime}\rangle_{S^{\prime}}\langle\eta_{+}|_{S}+\sum_{m\neq i^{\prime}}d_{k,m}|m\rangle_{S^{\prime}}\langle\eta_{-}|_{S}+\sum_{m\neq i^{\prime}}\sum_{l\neq i,j}f_{k,m,l}|m\rangle_{S^{\prime}}\langle l|_{S}.$ (S.127)
@@
$\langle i^{\prime}|_{S^{\prime}}\Lambda(\psi_{S})|i^{\prime}\rangle_{S^{\prime}}=0,\enskip\forall|\psi\rangle_{S}\mbox{ s.t. }\langle{\eta_{+}}|{\psi}\rangle_{S}=0.$ (S.128)